what percent of 62 should be added to 20% of 100 to give 92?
Answer #1

Let x percent of 62 should be added to 20% of 100 to give 92. i.e.

\frac{x}{100}\left ( 62 \right ) + \frac{20}{100}\left (100 \right ) =92

\frac{x}{50}\left (31 \right ) + 20=92

\frac{31x}{50} =92-20\\ \frac{31x}{50} =72\\ x= 72\left ( \frac{50}{31}\right )\\ x= \frac{3600}{31}

Which is required result.